OpenStudy (anonymous):

I am CONFIZZLED with this System of Equations Algebra I Question!!! Please help! Medal, Fan, Testimonial!

OpenStudy (anonymous):

A system of equations is shown below. {-3x + 7y = -16 {-9x + 5y = 16 Part A: Create an equivalent system of equations by replacing one equation with the sum of that equation and a multiple of the other. Show the steps to do this. (6 points) Part B: Show that the equivalent system has the same solution as the original system of equations. (4 points)

OpenStudy (anonymous):

A person told me to do this for Part A, is it correct?:

OpenStudy (anonymous):

Confizzled...lol. That is all.

OpenStudy (anonymous):

\[-3x+7y=-16\]\[\times3~~~~\times3~~~~\times3\]------------------\[-9x + 5y=16\] \[-9x+21y=-48\]\[+~9x+5y=16\]---------------\[0+26y=32\]

OpenStudy (potatoes.ramu):

tHAT'S WHAT i THOUGHT TOO

OpenStudy (potatoes.ramu):

Sorry that was in capitals my bad

OpenStudy (mathmale):

D&G: Please just multiply the first equation by -3. Then add the result to the second equation, OK? You'll understand this better if you do it yourself. Mind sharing an image of your results, when you 're done? solve the resulting system for y.

OpenStudy (anonymous):

Okay, I will do that now. @mathmale

OpenStudy (anonymous):

{-3x + 7y = -16 {-9x + 5y = 16 -3x + 7y = -16 *3 *3 *3 -------------- 9x + 21 = 48 + -9x + 5y = 16 ----------------- 0 + 26y = 64

OpenStudy (mathmale):

You're absolutely on the right track. But there are minor arith. mistakes to correct. First, you're right to multiply the first equation by 3, if you're going to change the signs later, but I did ask you to multiply the first equation by -3. Don't bother to do this over; stay with what you have. (I'm continuing to type to you.)

OpenStudy (anonymous):

omg @mathmale I did multiply by -3, I just forgot to put the sign here.

OpenStudy (mathmale):

One more thing: you should end up with 9x - 21y = 48. See the different sign in front of the y term? Review your work and determine what went wrong.

OpenStudy (anonymous):

Well, does a negative number multiplied by a negative number mean a negative product?

OpenStudy (mathmale):

Focus on the 5th equation that you 've typewritten to me.

OpenStudy (mathmale):

You should now have the following system: 9x - 21y = 48 -9x +5y = 16 Add the contents of each column together. You'll end up with what?

OpenStudy (anonymous):

I'll end up with "0 - 16y = 64"?

OpenStudy (mathmale):

Great. Bravo! And so, y=?

OpenStudy (anonymous):

y = -4

OpenStudy (mathmale):

Perfect. Now take that y=-4 and substitute it into either of the original two equations; then solve for x.

OpenStudy (anonymous):

Okay, thank you!

OpenStudy (mathmale):

Are you completely satisfied with the outcomes of our discussion, or is there anything else about it thatyou'd like to discuss?

OpenStudy (anonymous):

I am completely satisfied. ;]

OpenStudy (anonymous):

I might tag you if I need any help. I might need some more, but bot at the moment. Thank you.

OpenStudy (mathmale):

Wonderful working with you; hope to have the pleasure of doing so again. And warmest thanks for that very kind testimonial!! MM